OpenStudy (anonymous):

which of the following is not an identity? A. (tanx+cotx)^2=csc^2x+sec^2x B. 1+cos^2x/sin^2x=1 C. tan^2x=sec^2-sin^2-cos^2x D. (cosx+sinx)^2=1+2cosx sinx

OpenStudy (anonymous):

I know it's B but what about A?

OpenStudy (amoodarya):

\[(tanx+cotx)^2=?\csc^2x+\sec^2x\\tan^2x +\cot^x +2\tan x \cot x=\\tan^2x+\cot^2x +2=\\(\tan^2x+1)+(1+\cot^2x)=\\sec^2x+\csc^2x\]so A is true relation

OpenStudy (anonymous):

Thank you the answer B was the answer like you said